OpenStudy (anonymous):

p(x)=2x^4-6x+1 Is this an even or odd function or neither? I tried solving it and got p(-x)=2x+6x+1 but I'm not sure if it is odd or neither. Can someone explain this to me please?

ganeshie8 (ganeshie8):

you have all the data, just conclude :)

ganeshie8 (ganeshie8):

\(p(x)=2x^4-6x+1\) \(p(-x)=2x^4+6x+1 \)

ganeshie8 (ganeshie8):

Since \(p(x) \ne p(-x)\), its not an even function

ganeshie8 (ganeshie8):

Since \(p(-x) \ne - p(x)\), its not an odd function either

ganeshie8 (ganeshie8):

So... ? :)

OpenStudy (anonymous):

it's neither, thanks! :D

ganeshie8 (ganeshie8):

yu wlcme ! good job !!!

OpenStudy (anonymous):

Thanks : )Can you help me with another problem that's similar to this one?

ganeshie8 (ganeshie8):

shoot

OpenStudy (anonymous):

f(x)= 4^3sqrt -x ........ the 3 is in the tip of the sqrt I forget what that's called.

ganeshie8 (ganeshie8):

\(\huge f(x) = 4 \sqrt[3]{-x}\)

ganeshie8 (ganeshie8):

like that ?

ganeshie8 (ganeshie8):

or, \(\huge f(x) = 4 ^\sqrt[3]{-x}\) like this ?